OpenStudy (anonymous):

which of the following expressions is not equal to 1/15? a. 5^5 x 5^0 x 5 x 5^4 b. 5^2 x 5^-6 x 5^3 x 5^-1 c. 5 x 5^0 x 5^-1 x 5^-2 d. 5^6 x 5^-10 x 5^3 x 5^-1

OpenStudy (anonymous):

Are you sure that is isn't 1/25

OpenStudy (anonymous):

yeah it is sorry, my mistake :S

OpenStudy (anonymous):

Definately the first one doesn't equal that

OpenStudy (anonymous):

Only the first equation doesn't equal that

OpenStudy (anonymous):

@rld613: thanks, hey do you mind helping me on other questions? (:

OpenStudy (anonymous):

\[\{a. ,b. ,c. ,d.\}\to \left\{9765625,\frac{1}{25},\frac{1}{25},\frac{1}{25}\right\} \]